2. What is the pressure in atm of 1.35 x 102 mole of oxygen gas (O2) if it is compressed in a tube with 8000 mL at 20C?
1
Expert's answer
2020-05-16T13:42:48-0400

According to the ideal gas law:

pV = nRT,

where p - pressure, V - volume, n - number of moles, R - ideal gas constant, T - temperature.

From here:

p = nRT / V

As n = 1.35 × 102 mol, R = 0.08206 L atm / mol K, V = 8000 ml = 8 L, and T = 20 °C = 293.15 K:

p = 1.35 × 102 mol × 0.08206 L atm / mol K × 293.15 K / 8 L = 406 atm


Answer: 406 atm